25 years ago, Hailey was five more than one-third as old as Jason was. Today, Jason is twenty six less than two times the age of Hailey. How old is Jason?
How do I figure this out.

Answer by stanbon(75887)

25 years ago, Hailey was five more than one-third as old as Jason was. Today, Jason is twenty six less than two times the age of Hailey. How old is Jason?
---------------
Let Jason's age be J
Let Hailey's age be H
-----------------
Equations:
(H-25) = (1/3)(J-25)
J = 2H-26
------------------
Substitute for "J" and solve for "H".
---
H-25 = (1/3)(2H-26-25)
3H-75 = 2H-51
H = 24 (Hailey's age now)
----
Since J = 2H-26, J = 48-26 = 22 (Jason's age now)
=====================================================
cheers,
Stan H.